OpenStudy (anonymous):

Determine the inflection points of f(x)=x-lnx if any !!! HELP PLEASE THANK YOU

OpenStudy (anonymous):

this is what I got for my second derivative

OpenStudy (anonymous):

maybe it should be positive

OpenStudy (jhannybean):

\[\large f(x) = x-\ln(x)\]\[\large f'(x) = 1 - \frac{1}{x}\]\[\large f'(x)=1-1x^{-1}\]\[\large f''(x) = 0+(-1)(-1)x^{-2}\]\[\large f''(x)=\frac{1}{x^2}\]

OpenStudy (anonymous):

how can I set this to zero?

OpenStudy (jhannybean):

Now set \(\large \frac{1}{x^2}\) to 0 and solve for x to find inflection points. \[\large \frac{1}{x^2}=0\] Since you annot solve for this there is no solution. \[\large 1 = (0)(x^2) \]\[\large 1 \ne 0\]

OpenStudy (anonymous):

so does this mean no inflection points?

OpenStudy (jhannybean):

Mmhmm.

OpenStudy (anonymous):

thanks for the help!

OpenStudy (jhannybean):

No problem :) You can also check to see if there are inflection points by checking where the graph changes concavity. (i.e concave up \(\rightarrow\) concave down, or concave down \(\rightarrow\) concave up)